# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ 2.1 Cost-aware Portfolio Estimator
We assume the true weight vector $\bm{w}_{1}^{*}$ and weight difference vector $\bm{\delta}^{*}_{t}$ are sparse. Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$ , respectively. And the sparsity assumption means the cardinality $s_{t}=|\mathcal{A}_{t}|\ll p$ for $t=1,2,\ldots,m$. We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees. In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to obtain a sparse estimator of $\bm{w}_{1}^{*}$:
$\widehat{\bm{w}}_{1}:$ $=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+P_{\lambda}(\bm{w}),$ (2.3)
where $\widehat{\bm{\Sigma}}_{1}$ and $\widehat{\bm{\mu}}_{1}$ are based on samples $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, $d=n+1$, and $P_{\lambda}(\bm{w})=\sum_{j=1}^{p}P_{\lambda}\left(\bm{w}_{j}\right)$ is the penalty function. The oracle (when the true support set $\mathcal{A}_{1}$ is known) portfolio construction weight estimator is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle}}=\left(\widehat{\bm{\bm{w}}}_{\mathcal{A}_{1}}^{\text{oracle}},\mathbf{0}\right)=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
We assume the oracle solution is unique, namely,
$\nabla_{j}\mathcal{L}_{1,t}\left(\widehat{\bm{\bm{w}}}^{\text{oracle}}_{1}\right)+h^{\text{oracle}}=0\quad\forall j\in\mathcal{A}_{1},$ (2.4)
where $h^{\text{oracle}}=\frac{2-\gamma\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}_{\mathcal{A}_{1}}}$ is the Lagrange parameter and $\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})$.
The above discussion pertains to the construction of oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle}}$ during the first portfolio construction stage. For the reallocation stage, by replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the explanation extends naturally to the construction of oracle estimator $\widehat{\bm{\delta}}_{t}^{\text{oracle}}$ for the reallocation stage.
A popular choice of $P_{\lambda}(\cdot)$ is the Lasso (Tibshirani, 1996). However, the $\ell_{1}$ penalty is biased, requiring a strong Irrepresentable Condition to be selection consistent (Zhao and Yu, 2006). Nonconvex penalization method may outperform the $\ell_{1}$ method in terms of estimation accuracy and selection consistency. The nonconvex penalty function $P_{\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
@@
    $P_{\lambda}(t)$ is increasing and concave in $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
@@
    $P_{\lambda}(t)$ is differentiable in $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0):=P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
@@
    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in\left(0,a_{2}\lambda\right]$;
@@
    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with the pre-specified constant $a>a_{2}$, where $a_{1}$ and $a_{2}$ are two fixed positive constants.
The mild conditions on $P_{\lambda}(\cdot)$ are typically met by a wide range of regularizers, such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010). It is easy to verify that $a_{1}=a_{2}=1$ for the SCAD, and its derivative is defined as
$P^{{}^{\prime}}_{\lambda}(|\beta|)=\lambda\left\{I(|\beta|\leq\lambda)+\frac{(a\lambda-|\beta|)_{+}}{(a-1)\lambda}I(|\beta|>\lambda)\right\}$
for some $a>2$, and $P^{{}^{\prime}}_{\lambda}(0+)=\lambda$.
However, the nonconvex nature of the penalty term presents computational challenges due to the potential multiple local optima. To address this issue, Zou and Li (2008) proposed a feasible solution using LLA. This approach transforms the nonconvex penalized model into a sequence of weighted $\ell_{1}$ models, thereby leveraging the computational efficiency of the $\ell_{1}$ method. Zou and Li (2008) suggested using the properly initialized one-step LLA estimator instead of the computationally intensive fully converged estimator. It has been demonstrated that the one-step LLA estimator enjoys the strong oracle property when $\bm{X}^{\top}\bm{X}/n\rightarrow\mathbf{M}$ for some positive definition matrix $\mathbf{M}$. Subsequently, Fan et al. (2014) extended this study to high-dimensional settings and showed that initializing the LLA algorithm with a Lasso solution leads to an oracle solution with high probability. The computation algorithm for first portfolio construction is summarized in Algorithm 1. For the computation algorithm at reallocation stages, it suffices to replace $\bm{w}$ with $\bm{\delta}$ and substitute the subscript $1$ with $t$.
@@
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\prime}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\prime}$
2:for $k=1,2,...$ do repeat the LLA iteration till convergence
3:     Obtain $\widehat{\bm{w}}_{t}^{(k)}$ by solving the following optimization problem
$\widehat{\bm{w}}_{1}^{(k)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(k-1)}\cdot\left|w_{j}\right|$
4:     Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(k)}$ with $\widehat{\theta}_{j}^{(k)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(k)}\right|\right)$
@@
For $t=1,2,\ldots,m$, we assume that the estimators for mean and covariance matrix satisfy the following element-wise max norm deviation condition for some absolute constants $c_{1},c_{2},M_{1},M_{2}$:
$\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{1}}),$ (2.5)
$\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{2}})$
When the returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}$ are independent and identically distributed sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, condition 2.5 is satisfied by sample mean and covariance matrix (see Lemma S.1 in the supplementary materials). Moreover, in high-dimensional settings, theoretically refined estimators, such as linear and nonlinear shrinkage estimators, are expected to better satisfy the assumptions outlined above. Notice that we can relax condition 2.5 and allow dependent return data, including strong-mixing ($\alpha$-mixing), $m$-dependent sequence, and physical (or functional) dependence measure for casual time series (Chang et al., 2024). Nonetheless, we do not provide the full theory result of CAPE under dependent return data here for simplicity.
We further assume the estimator of covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the following restricted strong convexity (RSC) condition:
$\forall\mathbf{\Delta}\in\mathbb{R}^{p},\mathbf{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\mathbf{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1}~{}~{}\text{where}~{}~{}v>0,\tau\geq 0.$ (2.6)
The sample covariance matrix is typically semi-positive definite and $\forall\bm{\Delta}\in\mathbb{R}^{p},\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq 0$, so the RSC condition hold trivially for $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\geq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. We only require the condition to hold over a cone of the form $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\leq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. The RSC condition is widely adopted in high-dimensional estimation and is satisfied with high probability if the return vector is i.i.d. sub-Gaussian random variables (Loh and Wainwright, 2015). To establish the non-asymptotic error bounds, we impose the following assumptions:
@@
    The estimators $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ the estimators satisfy the element-wise max norm deviation condition (2.5) and the RSC condition (2.6);
@@
    There exists $C_{1},C_{2}>0$, s.t. $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}\|_{1}<C_{1}$, $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{2},\ \forall\bm{r}\in\{-1,0,1\}^{s_{t}}$;
@@
    The covariance matrix satisfies $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty$ and there exists $C_{3}>0$, such that $|||\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}<C_{3}$;
  • (A4)
    Suppose the theoretical minimizer $\bm{w}_{1}^{*}$ is sparse with cardinality $s_{1}=\|\bm{w}^{*}_{1}\|_{0}$ and the minimal signal strength of $\bm{w}_{1}^{*}$ satisfies that $\|\bm{\bm{w}}_{\mathcal{A}_{1}}^{*}\|_{\min}>(a+1)\lambda$.
Assumptions (A1)–(A3) hold for both the first portfolio construction stage and the reallocation stage, implying that $t=1,2,\ldots,m$. As demonstrated by Liu et al. (2024), the sample mean and covariance matrix satisfy the condition (2.5) in Assumption (A1), if the returns $\bm{R}_{d-n},\ldots,\bm{R}_{d-1}$ are i.i.d. sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$, and $\log p<n$. To improve the estimation accuracy, both linear and nonlinear shrinkage estimator employed in this study are assumed to satisfy condition (2.5). In Assumption (A2), for simplicity, we denote the submatrix of $\bm{\Sigma}_{t}$ consisting of the rows and columns indexed by $\mathcal{A}_{t}$ as $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$, and similarly, the subvector of $\bm{\mu}_{t}$ containing the elements indexed by $\mathcal{A}_{t}$ is represented as $\bm{\mu}_{\mathcal{A}_{t}}$. This notation will be adopted consistently in the subsequent discussion. Assumption (A2) is introduced to ensure that the $\ell_{1}$-norm of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ remain bounded, thereby preventing divergence. Similar assumptions can be found in Cai et al. (2022) and Liu et al. (2024). And the assumption constraining the $\ell_{1}$-norm of the parameter to prevent divergence also appears in Loh and Wainwright (2015). Furthermore, Assumption (A3) is less restrictive than the Irrepresentable Condition required for the Lasso, which restricts the strength of interactions between variables in $\mathcal{A}_{t}$ and $\mathcal{A}^{c}_{t}$. It ensures that the influence of variables outside $\mathcal{A}_{t}$ are not excessively large, preserving sparsity or stability in the covariance structure. Finally, the sparsity assumption in Assumption (A4) is reasonable and has also been employed in Du et al. (2022). Assumption (A4) also imposes the minimal signal strength condition, which is required for selection consistency, and the similar condition can be found in Fan et al. (2014), Loh and Wainwright (2017).
2.2 CAPE with Quadratic Transaction Cost

### v2

@@ 2.2 Cost-aware Portfolio Estimator
Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$, respectively. And we assume the cardinality $s_{t}=|\mathcal{A}_{t}|\leq s_{0}$, for $t=1,2,\ldots,m$. We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees.
To address the computational intractability of the $\ell_{0}$-constrained problems in (2.4) and (2.5), we approximate the $\ell_{0}$ constraint with a general sparsity-inducing penalty $p_{\lambda}(\cdot)$, yielding the following penalized formulations.
At the initial portfolio construction stage ($t=1$):
$\widehat{\bm{w}}_{1}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\,\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+\sum_{j=1}^{p}p_{\lambda}(|w_{j}|)\quad\text{s.t.}\quad\bm{w}^{\top}\mathbf{1}=1.$ (2.6)
At the reallocation stage ($t\geq 2$):
$\widehat{\bm{\delta}}_{t}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}-\gamma\,\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta})+\sum_{j=1}^{p}p_{\lambda}(|\delta_{j}|)\quad\text{s.t.}\quad\bm{\delta}^{\top}\mathbf{1}=0.$ (2.7)
Here, $p_{\lambda}(\cdot)$ denotes the sparsity-inducing penalty, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ are the estimators of the mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively, for $t=1,2,\dots,m$.
Two popular choices of $p_{\lambda}(\cdot)$ are the $\ell_{1}$ penalty (i.e., Lasso), and the Smoothly Clipped Absolute Deviation (SCAD) penalty. While the Lasso penalty is convex and yields efficient computation, it introduces bias and requires the strong Irrepresentable Condition to achieve variable selection consistency (Zhao and Yu, 2006). In contrast, nonconvex penalties like SCAD can reduce the estimation bias and achieve better selection consistency under weaker conditions.
The SCAD penalty (Fan and Li, 2001) satisfies the following regularity conditions:
@@
    $P_{\lambda}(t)$ is increasing and concave on $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
@@
    $P_{\lambda}$ is differentiable on $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
@@
    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in(0,a_{2}\lambda]$;
@@
    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with some constant $a>a_{2}$,
where $a_{1}$ and $a_{2}$ are fixed positive constants.
The SCAD penalty has $a_{1}=a_{2}=1$ and its derivative is given by
$P^{\prime}_{\lambda}(|\tau|)=\lambda\left\{I(|\tau|\leq\lambda)+\frac{(a\lambda-|\tau|)_{+}}{(a-1)\lambda}I(|\tau|>\lambda)\right\}$
for some $a>2$, and $P^{\prime}_{\lambda}(0+)=\lambda$.
Building on the general penalized formulations in (2.6) and (2.7), and leveraging the discussion of commonly used penalties such as Lasso and SCAD, we develop a novel cost-aware optimization framework specifically designed for high-dimensional portfolio selection. Specifically, we define the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) at the first portfolio construction stage as follows:
$\widehat{\bm{w}}_{1}=\mathop{\arg\min}_{\bm{w}_{1}\in\mathbb{R}^{p}}\;\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1})+P_{\lambda}(\bm{w}_{1})\quad\text{s.t.}\quad\bm{w}_{1}^{\top}\mathbf{1}=1,$ (2.8)
where $P_{\lambda}(\bm{w}_{1})=\sum_{j=1}^{p}P_{\lambda}(\bm{w}_{t,j})$ is SCAD penalty. This formulation simultaneously accounts for portfolio risk, expected return, transaction cost, and sparsity, making it particularly suitable for high-dimensional financial settings.
To facilitate theoretical analysis, we consider the oracle estimator which assumes knowledge of the true support set $\mathcal{A}_{1}$. The oracle portfolio weight estimator is defined by restricting the weights outside $\mathcal{A}_{1}$ to zero:
$\widehat{\bm{w}}_{1}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{w}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\,\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}).$
We assume the oracle solution satisfies the first-order optimality conditions:
$\nabla_{j}\mathcal{L}_{n,1}(\widehat{\bm{w}}_{1}^{\mathrm{oracle}})+h^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{1},$
where the Lagrange multiplier
$h^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}},$
$\mathbf{g}:=\partial\mathbf{C}_{1}(\bm{w}_{1})/\partial\bm{w}_{1}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{1}(\bm{w}_{1})$
and
$\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
The above describes the oracle estimator during the first initial portfolio construction stage. Similarly, for the reallocation stage with $t\geq 2$, the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) are defined as
$\widehat{\bm{\delta}}_{t}=\mathop{\arg\min}_{\bm{\delta}_{t}\in\mathbb{R}^{p}}\;\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t})+P_{\lambda}(\bm{\delta}_{t})\quad\text{s.t.}\quad\bm{\delta}^{\top}_{t}\mathbf{1}=0,$
where $P_{\lambda}(\bm{\delta}_{t})=\sum_{j=1}^{p}P_{\lambda}(\bm{\delta}_{t,j})$ is SCAD penalty and the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ denote estimators of the true mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively.
The corresponding oracle estimator is
$\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{\delta}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\,\bm{\delta}^{\top}_{t}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}),$
which satisfies
$\nabla_{j}\mathcal{L}_{n,t}(\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}})+h_{t}^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{t},$
where
$h_{t}^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{t}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}},$
$\mathbf{g}:=\partial\mathbf{C}_{t}(\bm{\delta}_{t})/\partial\bm{\delta}_{t}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{t}(\bm{\delta}_{t}),$
and
$\mathcal{L}_{n,t}(\bm{\delta}_{t})=\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}).$
However, the nonconvexity of SCAD poses significant computational challenges due to the presence of multiple local minima. To overcome this issue, Zou and Li (2008) proposed the local linear approximation (LLA) algorithm, which iteratively approximates the nonconvex problem by a sequence of weighted $\ell_{1}$-penalized problems, thereby inheriting the computational advantages of convex optimization. Rather than pursuing full convergence, they advocated for a properly initialized one-step LLA estimator, which is computationally efficient and shown to possess the strong oracle property under the condition that $\bm{X}^{\top}\bm{X}/n\to\mathbf{M}$ for some positive definite matrix $\mathbf{M}$. This framework was later extended by Fan et al. (2014) to high-dimensional settings, demonstrating that initializing the LLA algorithm with a Lasso solution yields an oracle estimator with high probability. The algorithmic procedures for the first portfolio construction and subsequent reallocation stages are presented in Algorithm 1 and Algorithm 2, respectively. Here, $\widehat{\bm{w}}_{1}^{\text{initial}}$ and $\widehat{\bm{\delta}}_{t}^{\text{initial}}$ denote initial estimators of $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$, respectively, satisfying certain regularity conditions. The specific definitions of these initial estimators, tailored to different transaction cost structures, are provided in Sections 2.3 and 2.4.
@@
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\top}$
2:for $l=1,2,$... do repeat the LLA iteration till convergence
3:  Obtain $\widehat{\bm{w}}_{t}^{(l)}$ by solving the following optimization problem
$\widehat{\bm{w}}_{1}^{(l)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|w_{j}\right|$
4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(l)}\right|\right)$
@@
Algorithm 2 Local Linear Approximation (LLA) Algorithm for Reallocation
1:Initialize $\widehat{\bm{\delta}}_{t}^{(0)}=\widehat{\bm{\delta}}_{t}^{\text{initial}}$ and compute the adaptive weight:
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,p}^{(0)}\right|\right)\right)^{\top}$
2:for $l=1,2,$... do repeat the LLA iteration till convergence
3:  Obtain $\widehat{\bm{\delta}}_{t}^{(l)}$ by solving the following optimization problem
$\widehat{\bm{\delta}}_{t}^{(l)}=\min_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\mathcal{L}_{n,t}(\bm{\bm{\delta}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|\delta_{j}\right|$
4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,j}^{(l)}\right|\right)$
5:end for
6:Output: The portfolio allocation $\widehat{\bm{\delta}}_{t}^{\text{LLA}}$
To establish the non-asymptotic error bounds, we impose the following assumptions:
@@
    There exist constants $C_{1},C_{2}>0$, for $t=1,2,\ldots,m$, such that
    $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{1},\quad\forall\bm{r}\in\{-1,0,1\}^{s_{t}},$
[…]
    $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty.$
@@
    For $t=1,2,\ldots,m$ and some absolute constants $c_{1},c_{2},M_{1},M_{2}$,
    $\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{1}}),$ (2.9)
    $\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{2}}).$
@@
    The estimated covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the restricted strong convexity (RSC) condition:
    $\forall\bm{\Delta}\in\mathbb{R}^{p},\quad\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1},$ (2.10)
    where $v>0$, $\tau\geq 0$, and $t=1,2,\ldots,m$.
Remark 1.
Assumption (A1) guarantees three key properties. The bound on $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}$ ensures that the $\ell_{1}$-norms of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ are controlled, preventing divergence in the estimation process. This condition aligns with assumptions in Cai et al. (2024). The constraint on $|\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}$, while weaker than the irrepresentable condition required for the Lasso, limits the correlation between active and inactive variables, promoting both sparsity and stability. Finally, the eigenvalue condition guarantees that $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ is well-conditioned, which facilitates stable matrix inversion and avoids ill-posedness when estimating coefficients restricted to the active set $\mathcal{A}_{t}$.
Assumption (A2) holds when the sample mean vector and sample covariance matrix are used as estimators, provided that the returns $\bm{R}_{(t-1)n+1},\ldots,\bm{R}_{tn}$ are i.i.d. sub-Gaussian vectors with mean $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, and that $\log p<n$, as shown in Liu et al. (2024). Furthermore, we show that the linear shrinkage estimator (Ledoit and Wolf, 2020) for the population covariance matrix also satisfies (A2) under the same sub-Gaussianity assumption and mild structural conditions on $\bm{\Sigma}_{t}$ in the supplementary material. The linear shrinkage estimator has more favorable properties and, in high-dimensional settings (e.g., $p$ comparable to or exceeding $n$), provides well-conditioned covariance estimates that make the classical Markowitz mean–variance model (2.1) and its transaction-cost extension (2.3) feasible and numerically stable. Moreover, the bounds in (2.9) can be extended to settings with dependent observations, including strong-mixing ($\alpha$-mixing) processes, $m$-dependence, and time series exhibiting physical (or functional) dependence (Chang et al., 2024). For simplicity, we focus on the i.i.d. case in this paper and defer the dependent setting to future work.
Assumption (A3) is a standard requirement in high-dimensional estimation. Loh and Wainwright (2015) established that the sample covariance matrix satisfies the RSC condition with high probability under sub-Gaussian assumptions. We further show that both linear and nonlinear shrinkage estimators (Ledoit and Wolf, 2020) satisfy (A3), as detailed in the supplementary material.
2.3 CAPE-S with Quadratic Transaction Cost